# Local Mechanisms of Compositional Generalization in Conditional Diffusion

### v2

@@ Abstract
Conditional diffusion models appear capable of compositional generalization, i.e., generating convincing samples for out-of-distribution combinations of conditioners, but the mechanisms underlying this ability remain unclear. To make this concrete, we study length generalization, the ability to generate images with more objects than seen during training. In a controlled CLEVR setting (Johnson et al., 2017), we find that length generalization is achievable in some cases but not others, suggesting that models only sometimes learn the underlying compositional structure. We then investigate locality as a structural mechanism for compositional generalization. Prior works proposed score locality as a mechanism for creativity in unconditional diffusion models (Kamb & Ganguli, 2024; Niedoba et al., 2024), but did not address flexible conditioning or compositional generalization. In this paper, we prove an exact equivalence between a specific compositional structure (conditional projective composition) (Bradley et al., 2025) and scores with sparse dependencies on both pixels and conditioners (local conditional scores). This theory also extends to feature-space compositionality. We validate our theory empirically: CLEVR models that succeed at length generalization exhibit local conditional scores, while those that fail do not. Furthermore, we show that a causal intervention explicitly enforcing local conditional scores restores length generalization in a previously failing model. Finally, we investigate feature-space compositionality in color-conditioned CLEVR, and find preliminary evidence of compositional structure in SDXL.
@@ 1 Introduction
Figure 1: Length generalization in location-conditioned CLEVR models. We study length generalization in location-conditioned models trained on images with 1-3 objects, and tested on 0-10 locations (0 and $\geq 4$ locations are out-of-distribution), with red dots indicating the conditioned locations at test-time. Experiments 1 and 2 have identical architectures and training data and differ only in the design of their conditioners. In Experiment 1, the conditioner labels the locations of all objects in the scene; the model successfully length-generalizes up to 9 locations (including 0). In Experiment 2, the conditioner labels the location of only a single object (randomly selected); the model largely fails to length-generalize. Experiment 3 attempts to “fix” Exp. 2 with a causal intervention: we modify the model architecture to explicitly enforce local conditional scores, use the same training data and conditioning as Exp. 2, and find that length generalization is restored.
Figure 2: Locality structures in location-conditioned CLEVR models (Experiments 1, 2, 3 of Figure 1). All models are conditioned on 4 locations (OOD). Each column represents a timestep $t$. Top row shows the predicted denoised images via learned scores. Lower rows (evaluated at two conditioned and one unconditioned location) show heatmaps of the pixel gradient magnitude (average absolute of Jacobian from one pixel to all other pixels), and the conditional gradient magnitude marked with $\times$ (with the “gradient” estimated via a finite difference of the score computed with and without each conditioner). Locality metrics are shown in Figure 10. All models show pixel-locality at low and high noise. The length-generalizing Experiment 1 model exhibits strong conditional-locality at high noise, while the non-length-generalizing Experiment 2 model critically lacks conditional-locality: the scores at conditioned locations either ignore all conditioners or depend on non-local conditioners. Experiment 3 model has architecturally-enforced local conditional scores, which restore length generalization. These experiments support the theoretical equivalence between CPC and LCS, and highlight the importance of conditional-locality at high noise.
@@
Conditional diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019; Song et al., 2020) appear to possess remarkable compositional generalization capabilities. For example, text-to-image models (Dhariwal & Nichol, 2021; Rombach et al., 2022; Ramesh et al., 2022) generate convincing images for prompts like “a photograph of a cat eating sushi with chopsticks” that were (probably) not seen during training. These models may generalize by composing known concepts (e.g. cat+sushi) in novel ways. However, the extent of generalization in large-scale models is unclear as their train sets are not publicly known (perhaps they have seen cats eating sushi). Further, despite recent progress (Okawa et al., 2024; Park et al., 2024; Sclocchi et al., 2025; Kadkhodaie et al., 2023; Favero et al., 2025; Chen et al., 2024; Wang et al., 2024; Lukoianov et al., 2025), the mechanisms underlying compositional generalization remain unclear.
@@
We first propose a concrete and controlled setting in which to study compositional generalization: length generalization in location-conditioned models trained on CLEVR Johnson et al. (2017), a synthetic dataset of objects with various locations, shapes, and colors. Length generalization refers to the ability to generate more objects than seen in training – e.g., can a location-conditioned model trained on 1-3 objects and tested on $K>3$ locations actually generate images with $K$ objects at the correct locations? Prior work demonstrated length generalization of explicit composition of multiple diffusion models via linear score combination Du et al. (2023); Liu et al. (2022); Bradley et al. (2025). In contrast, we study length generalization of a single conditional model. By training on multi-object samples, we hope that this model can learn the underlying compositional structure of the data, hence length-generalize. We find empirically that, depending on the conditioning specifics, these models sometimes length-generalize and sometimes do not.
@@
Next, we study local mechanisms for compositional generalization. We build primarily on two lines of prior work: one on local mechanisms for creativity, and another on compositionality in diffusion. First, Kamb & Ganguli (2024); Niedoba et al. (2024) recently proposed that models learn local score functions, enabling creativity via mosaicing of local patches from different images. These works only study unconditional and class-conditional diffusion, however, and do not consider flexible conditioners such as those used in text-to-image diffusion, which are central to questions of compositional generalization. Also, they empirically observe delocalization in trained score networks at high noise, when global scene structure is often resolved (Wang & Vastola, 2023). It therefore remains unclear whether local mechanisms are relevant to compositional generalization in conditional diffusion models. Second, Bradley et al. (2025) propose a formal definition, called projective composition, of “correct” composition of multiple distributions Du et al. (2023); Liu et al. (2022). In this paper, we specialize projective composition to a single conditional distribution to provide a precise definition of composition structure.
@@
We develop a theoretical framework connecting compositional generalization with local mechanisms. Specifically, we generalize the concept of local scores to define local conditional scores: scores with sparse dependencies on both pixels and conditioners. That is, the score at a given pixel depends only on a subset of other pixels (such as a local neighborhood) and on only one or a few relevant conditioners (e.g. in the case of location-conditioning, only conditioners near the current pixel). We specialize projective composition (Bradley et al., 2025) to define a conditional projective composition – a conditional distribution that is a projective composition of its own individual conditionals. We then prove an equivalence between conditional projective composition and local conditional scores at all noise levels. We extend this theoretical framework to relate compositional structure and sparse score dependencies in feature-space (intuitively, if concepts like style+content compose in an appropriate disentangled feature space, then the score of each ‘style feature’ depends only on a sparse set of conditioners and other features related to style).
@@
We validate this theory through experiments. Returning to our location-conditioned CLEVR setting and comparing a model that we found to length-generalize with one that did not, we find that both models maintain pixel-locality (even at high noise levels, in contrast to CIFAR-10 and other models studied by Kamb & Ganguli (2024); Niedoba et al. (2024)). However, the length-generalizing model has sparse dependencies on its location-conditioners while the non-length-generalizing model does not. Further, we perform a direct causal intervention to test local conditional scores as a possible mechanism for composition generalization: we show that explicitly enforcing a local architecture restores length generalization in the model that previously failed. Finally, we investigate feature-space compositionality in color-conditioned CLEVR, and show preliminary SDXL experiments to explore compositional structure in real-world text-to-image models.
@@ 2 Length generalization in CLEVR
Train data Exp. 1 Exp. 2 Exp. 3 Color
1 object 1 1 6 1
1-2 objects 5 1+(0-1)
1-3 objects 9 1+(0-2) 9+(0-2) 4
1-4 objects 10 1+(0-3)
1-5 objects 10 1+(0-4) 10+(0-2) 7
1-6 objects 11 1+(0-5)
Table 1: Limits of length generalization in location- and color-conditioned CLEVR. The table lists the maximum value, $K_{\max}$, such that the model “succeeds” for every $1\leq K\leq K_{\max}$ (Section E.2). Exp. 1,2,3 refer to the location-conditioning experiments in Figure 1, Color refers to color-conditioning as in Figure 5. Parentheses e.g. +(0-2) indicate extra objects; untested configurations are left blank.
In this section we study length generalization in conditional diffusion models trained on CLEVR datasets Johnson et al. (2017), using a standard EDM2 U-net architecture Karras et al. (2022) (details in Section E.1). Figure 1 shows length generalization or lack thereof in two location-conditioned models trained on CLEVR images with 1-3 objects. In Experiment 1, the location-conditioning labels all objects in the scene, using a 2d integer array representing a 2d grid over the image via the count of objects whose center falls within the grid cell (typically zero or one), as shown in Figure 9. We find that this model successfully length-generalizes up to 9 objects, indicating that it has learned the underlying compositional structure of the data. In Experiment 2, however, length generalization fails. The setup is identical to that of Experiment 1 except for one difference: the conditioning only labels the location of a single randomly-selected object. (By analogy to text-to-image conditioning, this difference is like text describing everything in each image vs. describing a single object in each image.) Since Experiment 2 conditions on only one location during training, there are at least two strategies that could perfectly fit the training data: draw an object at every conditioned location (compositional), or draw an object at any conditioned location (non-compositional). However, these strategies are not equivalent at inference time, when we condition on more than one location. The model evidently chooses a non-compositional strategy that fits the training data but does not length-generalize. Experiment 3 “fixes” Experiment 2 via a local causal intervention motivated by our theory, as we discuss in Section 4 (note that since this model was trained on 1-3 object with only one location labeled, the expected inference-time behavior is to draw an object at each conditioned location, plus 0-2 extra objects). We also explore color-conditioning in CLEVR in Figure 5, which we expect to have a compositional structure in an appropriate feature-space, discussed further in Section 4. Table 1 gives a quantitative analysis of the limits of length generalization of the various location- and color-conditioned models, trained on $1$ to $M$ objects for $M=1,\ldots,6$ (Exp. 1 samples in Figure 8). The table reports an approximate measure of the maximum number of objects to which each model can consistently length-generalize (details in Section E.2).
@@ 3 Theory: compositionality and locality
Figure 3: Conditional projective composition (CPC) and local conditional scores (LCS). A CPC is a conditional distribution over a set of conditions $c_{\mathcal{J}}$ that factorizes independently into the marginals over $x_{M_{j}}$ conditioned on $c_{j}$, where $M_{j}$ are disjoint subsets. An LCS is a conditional score over a set of conditions $c_{\mathcal{J}}$ such that the score at each pixel $i$ depends only on a subset $N_{i}$ of other pixels (often a local neighborhood) as well as a subset $L_{i}\subset\mathcal{J}$ of conditions (for location-conditioning, often nearby conditioners). For certain choices of subsets, CPC and LCS are equivalent.
@@
In this section we present theory connecting compositional generalization to generalized local mechanisms. We first define local conditional scores (LCS), an extension of local scores Kamb & Ganguli (2024); Niedoba et al. (2024) to account for flexible conditioners. An LCS evaluated at a given pixel has sparse dependencies on other pixels (generalizing local neighborhoods) and sparse dependencies on conditioners. Next, we define conditional projective composition (CPC), a special case of projective composition (PC) proposed in Bradley et al. (2025) applied to a single conditional distribution. A conditional distribution that satisfies CPC factorizes into independent distributions over disjoint subsets of pixels that depend on a single condition. LCS and CPC are illustrated in Figure 3. We then prove that the score of a CPC is exactly an LCS: intuitively, compositional distributions have local scores. This result can be partially relaxed in an approximation that improves at higher noise, which could allow approximately-compositional structure to be resolved early in denoising. The theory can also be extended to feature space, to connect compositional structure (e.g. style+content) with sparse score dependencies (e.g. scores of ‘style features’ only depend on select conditioners and features relevant to style). The remainder of this section makes these claims precise.
@@
We generalize the idea of local scores (Kamb & Ganguli, 2024; Niedoba et al., 2024) to account for flexible compositional conditioners, [^1] Our definition breaks slightly from the originals in defining local scores in terms of a distribution rather than a finite training set, and also omits equivariance, which is unnecessary for our theory. which are central to compositional generalization. Let $p(x|c)$ be the true distribution over data $x\in\mathbb{R}^{n}$ conditioned on $c$. Let $x_{N}$ denote the restriction of $x$ to a subset of indices $N\subseteq[n]$, and $p(x_{N}|c)$ denote the marginal distribution of $p(\cdot|c)$ on $x_{N}$. Conditioners are represented as $c_{\mathcal{J}}=\{c_{j},j\in\mathcal{J}\},$ where $\mathcal{J}\subseteq\mathcal{J}_{\text{all}}$ is a subset of all possible conditioners. We assume that $p$ is defined for any combination of conditioners, even those not seen during training. A local conditional score at pixel $i$, $s^{t}[x|c_{\mathcal{J}}](i)$, depends on two subsets (Figure 3): $N_{i}$, a subset of pixels relevant to pixel $i$, and $L_{i}(\mathcal{J})$, a subset of conditions in $\mathcal{J}$ relevant to pixel $i$. In general, the subsets $N_{i}$ and $L_{i}$ need not be disjoint (although this will later be necessary to achieve CPC equivalence), and may contain multiple objects or conditioners.
@@
$s^{t}[x|c_{\mathcal{J}}](i)$ $:=\nabla\log p^{t}[x_{N_{i}^{t}}|c_{L_{i}^{t}(\mathcal{J})}](i),\quad\text{for all pixels $i$}.$ (1)
@@
To define a compositional structure for conditional distributions, we introduce conditional projective composition (CPC). We do not claim all distributions have this structure, but we will show that those that do also have a local score structure (LCS), suggesting a mechanism for compositional generalization. We specialize Bradley et al. (2025)’s pixel-space projective composition to the case where the concept distributions $p_{j}$ represent the a single conditional distribution $p(x|c_{j})$ conditioned on different $c_{j}$, and the background distribution $p_{b}(x)$ is $p(x|\emptyset)$ (i.e., with no conditioners active).
@@
Figure 4: Local causal intervention enables length generalization in a location-conditioned model trained on a single object. We test length generalization in two different location-conditioned models, both trained on images with only a single object (and conditioned on its single location). Left: A model with the standard EDM2 architecture of Experiment 1 does not length-generalize: it always generates exactly one object (even when conditioned on zero locations). Right: A model with the explicitly-enforced local architecture of Experiment 3 (see Section E.3) length-generalizes up to 6 objects, albeit with some artifacts (objects “merging” into each other). Although it does not perform as well as the Exp. 3 local model trained on 1-3 objects (Figure 1), any length-generalization after training on only one object is remarkable. (In Section E.2, we hypothesize that training on more objects, e.g. 1-3, may improve length-generalization by allowing the model to learn clusters of objects).

### v3

@@ Abstract
Conditional diffusion models appear capable of compositional generalization, i.e., generating convincing samples for out-of-distribution combinations of conditioners, but the mechanisms underlying this ability remain unclear. To make this concrete, we study length generalization, the ability to generate images with more objects than seen during training. In a controlled CLEVR setting (Johnson et al., 2017), we find that length generalization is achievable in some cases but not others, suggesting that models only sometimes learn the underlying compositional structure. We then investigate locality as a structural mechanism for compositional generalization. Prior works proposed score locality as a mechanism for creativity in unconditional diffusion models (Kamb and Ganguli, 2024; Niedoba et al., 2024), but did not address flexible conditioning or compositional generalization. In this paper, we prove an exact equivalence between a specific compositional structure (conditional projective composition) (Bradley et al., 2025) and scores with sparse dependencies on both pixels and conditioners (local conditional scores). This theory also extends to compositions of concepts (such as style+content) in feature-space. We validate our theory empirically: CLEVR models that succeed at length generalization exhibit local conditional scores, while those that fail do not. Furthermore, we show that a causal intervention explicitly enforcing local conditional scores enables length generalization in a previously failing model. Finally, we investigate SDXL and find that in pixel-space, spatial locality is present but conditional-locality is mostly absent; however, we find quantitative evidence of local conditional scores in the network’s learned feature-space.
@@ 1 Introduction
Figure 1: Length generalization in location-conditioned CLEVR models. We study length generalization in location-conditioned models trained on images with 1-3 objects, and tested on 1, 3, 6, 9 locations (6, 9 are OOD), with red dots indicating the conditioned locations at test-time. For each experiment, the rows correspond to different conditioners (1, 3, 6, or 9 locations) and the columns show 4 different samples. All models have the same architectures and training data and differ only in the design of their conditioners (see Figure 6). In Experiment 1, a grid-style conditioner labels the locations of all objects in the scene; the model successfully length-generalizes up to 7 locations. In Experiment 2, a grid-style conditioner labels the location of only a single object (randomly selected); the model fails to length-generalize (in this case, even 3 locations is OOD). In Experiment 3, a list-style conditioner labels the locations of all objects; this model fails to length-generalize beyond 3 objects. Additional samples shown in Figure 7. Experiment 2L applies a causal intervention to the failing Exp. 2: we modify the model architecture to explicitly enforce local conditional scores, use the same training data and conditioning as Exp. 2, and find that Exp. 2L length-generalizes while Exp. 2 failed (see also Section E.4 for an analogous Exp. 3L).
@@
Conditional diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019; Song et al., 2020) appear to possess remarkable compositional generalization capabilities. For example, text-to-image models (Dhariwal and Nichol, 2021; Rombach et al., 2022; Ramesh et al., 2022) generate convincing images for prompts like “a photograph of a cat eating sushi with chopsticks” that were (probably) not seen during training. These models may generalize by composing known concepts (e.g. cat+sushi) in novel ways. However, the extent of generalization in large-scale models is unclear as their train sets are not publicly known (perhaps they have seen cats eating sushi). Further, despite recent progress (Okawa et al., 2024; Park et al., 2024; Sclocchi et al., 2025; Kadkhodaie et al., 2023; Favero et al., 2025; Chen et al., 2024; Wang et al., 2024; Lukoianov et al., 2025), the mechanisms underlying compositional generalization remain unclear.
@@
We first propose a concrete and controlled setting in which to study compositional generalization: length generalization in location-conditioned models trained on CLEVR Johnson et al. (2017), a synthetic dataset of objects with various locations, shapes, and colors. Length generalization refers to the ability to generate more objects than seen in training – e.g., can a location-conditioned model trained on 1-3 objects and tested on $K>3$ locations actually generate images with $K$ objects at the correct locations? Prior work demonstrated length generalization of explicit composition of multiple diffusion models via linear score combination Du et al. (2023); Liu et al. (2022); Bradley et al. (2025). In contrast, we study length generalization of a single conditional model. By training on multi-object samples, we hope that this model can learn the underlying compositional structure of the data, hence length-generalize. We find empirically that, depending on conditioning and architecture specifics, these models sometimes length-generalize and sometimes do not.
@@
Next, we study local mechanisms for compositional generalization. We build primarily on two lines of prior work: one on local mechanisms for creativity, and another on compositionality in diffusion. First, Kamb and Ganguli (2024); Niedoba et al. (2024) recently proposed that models learn local score functions, enabling creativity via mosaicing of local patches from different images. These works only study unconditional and class-conditional diffusion, however, and do not consider flexible conditioners such as those used in text-to-image diffusion, which are central to questions of compositional generalization. It therefore remains unclear whether local mechanisms are relevant to compositional generalization in conditional diffusion models. Second, Bradley et al. (2025) propose a formal definition, called projective composition, of “correct” composition of multiple distributions Du et al. (2023); Liu et al. (2022). In this paper, we specialize projective composition to a single conditional distribution to provide a precise definition of composition structure.
@@
We develop a theoretical framework connecting compositional generalization with local mechanisms. Specifically, we generalize the concept of local scores to define local conditional scores (LCS): scores with sparse dependencies on both pixels and conditioners. That is, the score at a given pixel depends only on a subset of other pixels (such as a local neighborhood) and on only one or a few relevant conditioners (e.g. in the case of location-conditioning, only conditioners near the current pixel). We specialize projective composition (Bradley et al., 2025) to define a conditional projective composition (CPC) – a conditional distribution that is a projective composition of its own individual conditionals. We then prove an equivalence between conditional projective composition and local conditional scores at all noise levels. We extend this theoretical framework to relate compositional structure and sparse score dependencies in feature-space (intuitively, concepts like style+content will compose in feature-space if the score of each ‘style feature’ depends only on a sparse set of style-related conditioners and other features).
@@
We validate this theory through experiments. Returning to our location-conditioned CLEVR setting and comparing a model that we found to length-generalize with others that did not, we find that the length-generalizing model maintains pixel- and conditional-locality, while the non-length-generalizing models exhibit non-locality. We find that the correlation between length-generalization and conditional-locality also holds over a wider range of models with varying length-generalization. Further, we perform a direct causal intervention to test local conditional scores as a possible mechanism for composition generalization: we show that explicitly enforcing a local architecture enables length generalization in a model that previously failed. Finally, we present quantitative evidence of feature-space disentanglement in SDXL, connecting compositional structure with real-world text-to-image models.
Figure 2: (Left) Locality in location-conditioned CLEVR models. For Experiments 1, 2 and 3 of Figure 1 each conditioned on four locations, we visualize pixel-locality via heatmaps, and conditional locality via the intensity of the $\times$ marker, centered at a pixel in the lower left, over a range of timesteps. (Appendix F describes the locality measurements; Figure 12 plots locality metrics; Figure 9 shows more pixel locations.) The length-generalizing Exp. 1 model exhibits strong pixel- and conditional-locality, while the non-length-generalizing Exp. 2 and 3 models both lack conditional-locality (the scores depend on non-local conditioners); Exp. 3 also lacks pixel-locality. These experiments support the theoretical equivalence between CPC and LCS. (Right) Length generalization vs. conditioner locality for several models (different colors), each checkpointed early, mid, and late in training (different shapes). Details are in Section E.2.1. Length generalization and conditional locality are strongly correlated, and can emerge together over the course of training (e.g. orange, green, red models). Here, length-generalization ($x$-axis) is the number of locations to which the model can generalize ($K_{\text{max}}$ in Table 1) minus the maximum number on which it was trained (e.g. +4 for a model trained on 1-3 locations that generalizes to 7). The conditional locality ($y$-axis) metric is described in Appendix F.
@@ 2 Length generalization in CLEVR
Train data Exp.1 Exp.2 Exp.3 Exp.2L Exp.3L
1 object 1 1 1 5
1-3 objects 7 2 3 9 9
1-5 objects 10 3 6 10
Table 1: Upper limits of length generalization in location-conditioned CLEVR. The table lists the maximum value, $K_{\max}$, such that the model “sometimes succeeds” for every $1\leq K\leq K_{\max}$, as described in Section E.2. $K_{\max}$ is evaluated using an automated ResNet-18 classifier over 1024 samples per configuration. Results are shown for Exp. 1, 2, 3, 2L, 3L of Figure 1. Dashes indicate untested configurations.
In this section we study length generalization in conditional diffusion models trained on CLEVR datasets Johnson et al. (2017), using a standard EDM2 U-net architecture Karras et al. (2022) (details in Section E.1). Figure 1 shows length generalization or lack thereof in three location-conditioned models trained on CLEVR images with 1-3 objects. In Experiment 1, the location-conditioning labels all objects in the scene, using a 2d integer array representing a 2d grid over the image via the count of objects whose center falls within the grid cell (typically zero or one), as shown in Figure 6. We find that this model length-generalizes up to 7 objects. In Experiment 2, the setup is identical to that of Experiment 1 except that the conditioning only labels the location of a single randomly-selected object. Unsurprisingly, this model fails to meaningfully length-generalize beyond one conditioned location (note that when trained on 1-3 objects it may generate up to 2 “extra” objects at non-conditioned locations). Experiment 3 conditions on the 2D locations of all objects using a list-style conditioner which places the (embedded) xy-locations of each object in an array padded with enough slots for up to 10 objects, with each location placed in a randomly chosen slot. This model fails to length-generalize beyond the 3 locations it was trained on. A priori, we should not “expect” length-generalization in any of these models. Although the data is naturally compositional, there is no guarantee that a model will learn a compositional structure from examples with only 1-3 objects (which it could fit in many different ways). Evidently, the Experiment 1 model learns the underlying compositional structure of the data while the two other models do not. Table 1 gives a quantitative analysis of the limits of length generalization of the various location-conditioned models, trained on 1 object, 1-3 objects, or 1-5 objects, with samples shown in Figure 10. The table reports an approximate measure, $K_{\max}$, of the maximum number of objects to which each model can consistently length-generalize (details in Section E.2).
@@ 3 Theory: compositionality and locality
Figure 3: (Left) Conditional projective composition (CPC) and local conditional scores (LCS). A CPC is a conditional distribution over a set of conditions $c_{\mathcal{J}}$ that factorizes independently into the marginals over $x_{M_{j}}$ conditioned on $c_{j}$, where $M_{j}$ are disjoint subsets. An LCS is a conditional score over a set of conditions $c_{\mathcal{J}}$ such that the score at each pixel $i$ depends only on a subset $N_{i}$ of other pixels (often a local neighborhood) as well as a subset $L_{i}\subset\mathcal{J}$ of conditions (for location-conditioning, often nearby conditioners). For certain choices of subsets, CPC and LCS are equivalent.
@@
In this section we present theory connecting compositional generalization to generalized local mechanisms. We first define local conditional scores (LCS), an extension of local scores Kamb and Ganguli (2024); Niedoba et al. (2024) to account for flexible conditioners. An LCS evaluated at a given pixel has sparse dependencies on other pixels (generalizing local neighborhoods) and sparse dependencies on conditioners. Next, we define conditional projective composition (CPC), a special case of projective composition (PC) proposed in Bradley et al. (2025) applied to a single conditional distribution. A conditional distribution that satisfies CPC factorizes into independent distributions over disjoint subsets of pixels that depend on a single condition. LCS and CPC are illustrated in Figure 3. We then prove that the score of a CPC is exactly an LCS: intuitively, compositional distributions have local scores. We verify this empirically in Figure 2, discussed further in Section 4. This result can be partially relaxed in an approximation that improves at higher noise, which could allow approximately-compositional structure to be resolved early in denoising. The theory can also be extended to feature space, to connect compositional structure (e.g. style+content) with sparse score dependencies (e.g. scores of ‘style features’ only depend on select conditioners and features relevant to style). The remainder of this section makes these claims precise.
Background
Our theory builds upon two lines of prior work: local scores and projective composition. Kamb and Ganguli (2024); Niedoba et al. (2024) propose that diffusion models learn local score functions, enabling them to generate samples not in the training set through mosaicing of local patches. Both works define local scores essentially as follows: let $x_{N}$ denote the restriction of $x$ to a subset of indices $N\subseteq[n]$, and $p(x_{N})$ denote the marginal distribution of $p(x)$ on $x_{N}$. Intuitively, a local score at pixel $i$ depends only on a neighborhood of pixels centered at pixel $i$. That is, $s^{t}$ is a local score at time $t$ with neighborhood subsets $N_{i}^{t}$ (which may depend on the time $t$), if $s^{t}[x](i):=\nabla\log p^{t}[x_{N_{i}^{t}}](i),$ for all pixels $i$. However, these works study unconditional and class-conditional diffusion, not flexible conditioners central to compositional generalization. In this work, we will extend this concept to incorporate flexible conditioners and study compositionality.
Bradley et al. (2025) introduce projective composition (PC) as a formal definition for “correctly composing” multiple distributions $\{p_{b},p_{1},p_{2},\dots\}$, where $p_{b}$ is a “background” distribution and the $p_{i}$ are “concept distributions”. One possible construction of a projective composition is given by $p_{\mathcal{J}}(x):=p_{b}(x_{b})\prod_{j\in\mathcal{J}}p_{j}(x_{M_{j}}),$ where the $M_{j}$ are disjoint subsets corresponding to the $p_{j}$, respectively. In this paper, we will specialize PC to define a specific compositional structure for conditional distributions. However, in contrast to Bradley et al. (2025)’s focus on explicit compositions, our goal is to connect compositional structure with score locality.
Next, we present new definitions and theory that build on local scores and projective composition.
@@
We generalize the idea of local scores (Kamb and Ganguli, 2024; Niedoba et al., 2024) to account for flexible compositional conditioners, [^1] Our definition breaks slightly from the originals in defining local scores via a distribution rather than finite training set, and also omits equivariance, which is unnecessary for our theory. which are central to compositional generalization. Let $p(x|c)$ be the true distribution over data $x\in\mathbb{R}^{n}$ conditioned on $c$. Let $x_{N}$ denote the restriction of $x$ to a subset of indices $N\subseteq[n]$, and $p(x_{N}|c)$ denote the marginal distribution of $p(\cdot|c)$ on $x_{N}$. Conditioners are represented as $c_{\mathcal{J}}=\{c_{j},j\in\mathcal{J}\},$ where $\mathcal{J}\subseteq\mathcal{J}_{\text{all}}$ is a subset of all possible conditioners. We assume that $p$ is defined for any combination of conditioners, even those not seen during training. A local conditional score at pixel $i$, $s^{t}[x|c_{\mathcal{J}}](i)$, depends on two subsets (Figure 3): $N_{i}$, a subset of pixels relevant to pixel $i$, and $L_{i}(\mathcal{J})$, a subset of conditions in $\mathcal{J}$ relevant to pixel $i$. In general, the subsets $N_{i}$ and $L_{i}$ need not be disjoint (although this will later be necessary to achieve CPC equivalence), and may contain multiple objects or conditioners.
@@
$s^{t}[x|c_{\mathcal{J}}](i)$ $:=\nabla\log p^{t}[x_{N_{i}^{t}}|c_{L_{i}^{t}(\mathcal{J})}](i),\quad\forall i.$ (1)
@@
To define a compositional structure for conditional distributions, we introduce conditional projective composition (CPC). We do not claim all distributions have this structure, but we will show that those that do also have a local score structure (LCS), suggesting a mechanism for compositional generalization. We specialize Bradley et al. (2025)’s pixel-space projective composition to the case where the concept distributions $p_{j}$ represent a single conditional distribution $p(x|c_{j})$ conditioned on different $c_{j}$, and the background distribution $p_{b}(x)$ is $p(x|\emptyset)$ (i.e., with no conditioners active).